OpenStudy (anonymous):

I dont understand domain and range. y = –3(x – 10)2 + 5

OpenStudy (anonymous):

y = –3(x – 10)^2 + 5

OpenStudy (anonymous):



OpenStudy (campbell_st):

the domain is the x values that can be used...or are there any x values that makes the relationship undefined. e.g. y = 1/x in this case x cannot = 0 as 1/0 doesn't exist. or \[y = \sqrt{9 - x}\] x has to be greater less than or equal to 9... values higher than 9 will result in having to take a negative square root. the range is the y values generated by the x values. in your question the domain is all real x...... or you can put any real number into the relationship.. the range is\[y \le\] the largest output value is y = 5 when x = 0... for all other values of x the result will be less than 5 . (x - 10)^2 will always give positive values -3(x-10)^2 will result in negative values

OpenStudy (anonymous):

Thank You!